Status Report

NASA Space Station On-Orbit Status 24 Jun 2003

By SpaceRef Editor
June 24, 2003
Filed under , , ,
NASA Space Station On-Orbit Status 24 Jun 2003

All ISS systems continue to function nominally, except as noted previously or below.  This is Week 8 of Increment 7.

After breakfast at 2:40am EDT, work plan review, and morning DPC (daily planning conference) via S-band, CDR Yuri Malenchenko started the day with the TV downlink of the geophysical GFI-1 Relaksatsiya (“relaxation”) video footage which had not been completed satisfactorily on 5/20.  [The downlink of the recorded footage was initiated by the Russian SPP automated daily timeline system during a Russian comm pass on Daily Orbit 3 (DO3).  The videocassette of the experiment was then stowed.]

In a 1.5-hr. maintenance task in the FGB (funktsionalnyi-grusovoi blok), after the FGB cooling system was switched to Internal Loop #1, Malenchenko removed the Loop #2 (VGK2-1) replaceable pump panel (SPN) and installed  a new SPN unit delivered on 12P/Progress-259.

FE/SO Ed Lu meanwhile worked on the CSA-CP (compound specific analyzer-combustion products) system in the Lab, resupplying the equipment and its stowage kit with two new CSA-CP units with fresh carbon monoxide (CO) and hydrogen chloride (HCL) sensors, ten new battery packs, one new zero filter and two spare pump filters.  He then deployed the new CSA-CPs outside the stowage kit in the open environment and readied the old CSA-CP for subsequent data takes.  [CSA-CP is a toxicology monitoring device for early CO detection/warning and post-fire decontamination.  It also checks for HCl and other combustion products and contaminants.]

Next, Lu conducted a “zero” calibration run on the CSA-CP, then activated the device for 5-min. readings in SM and Lab before deactivating it and attaching it at its regular location at the SM port wall central post.

A planned 1-hr. activity by Yuri Malenchenko to reconfigure electronic cables in the SM to connect the Regul-OS antenna feeder unit (AFU) to the antenna feed in a new arrangement was deferred.

The Science Officer had 2.5 hours for performing the scheduled functional video downlink test of the HRF (human research facility) Ultrasound equipment (UVND).  [Ed Lu’s task had two main objectives:  First, setting up and activating the HRF Ultrasound (U/S), followed by downlinking a series of video test pattern images to the ground so that HRF specialists can assess real-time video image downlink quality from the U/S.  Then, in an experiment of remote-guidance telemedicine with privacy-protected access, Space Medicine specialists at the JSC TSC (Telescience Support Center) were to guide him via privatized audio channel through acquisition of a series of human subject internal images, in order to assess the degree of downlinked U/S video and still-image degradation from clinical perspective.  The U/S equipment is part of HRF Rack 1, launched on 5A.1 in March 2001. Its applications include echocardiography, abdominal U/S (deep organ), vascular U/S, muscle and tendon U/S, transcranial U/S and U/S contrast studies.  The procedure was jointly prepared and conducted by the JSC TSC, POC (Huntsville Payload Operations Center), and MCC-H.]

In the SM, Yuri Malenchenko performed a 2-hr. hardware upgrade on the BRPK condensate separation and pumping units by replacing the old pipe conduits, including safety valves, with new units.  [The latter have no filters and use higher safety valve trip pressures.  Purpose of the upgrade is to prevent unwanted flow of the fluid through the bypass line during condensate transfer cycles.]

Yuri also completed the regular weekly checkout of the IP-1 air flow sensors in the various RS hatchways and the passage between FGB and Node.

Ed Lu completed the periodic transfer of accumulated physical exercise data files of the TVIS treadmill and RED resistive exerciser to the MEC (medical equipment computer), for subsequent downlink to the ground.  [The downloads are scheduled after more than four sessions have been performed since the last session.]

Ed also conducted the daily routine maintenance of the SM’s SOZh environment control & life support system, while Yuri later prepared the daily “delta” file for updating the inventory management system (IMS).

At 11:50am EDT, both crewmembers participated in an interactive televised PAO interview (Ku-band & S-band) with’s Steven Young.  [The interview was taped for later placement on the client’s website, and also broadcast live on NASA-TV.  It focused primarily on how the crew is spending its time onboard; the arrival of Progress 11 and its deliveries; daily life aboard ISS; ability to maintain the station with only two persons; thoughts on the recovery from the Columbia accident, the future of international cooperation for further development of ISS, and space tourism.  Questions had also been submitted by visitors to the website.]

The crew completed their regular daily 2.5-h program of physical exercise, on TVIS treadmill, RED expander and, for Yuri, on the Russian VELO cycle ergometer with load trainer.

Today’s CEO (crew earth observation) targets, limited in the current XPOP attitude by flight rule constraints on the use of the Lab nadir/science window, and including the targets of the Lewis & Clark 200-year memorial locations, were Tucson forest fire, AZ (Dynamic event:  Many dozens of homes feared destroyed in the Catalina Mts. north of Tucson as the nation’s largest fire rages.  Looking right for the smoke clouds moving northward.  Nadir view of Phoenix), Denver, Colorado (nadir view of the city  detectable immediately east of the dark forested Front Range of the Rockies), Taiwan smog (looking right of track for probable smog buildup on the heavily industrialized west side of the island.  The west coast is often invisible from space due to industrial haze.  A second series of images, slightly overexposed, might reveal structure in the smog mass [structure indicates varying density and transport direction]), Angolan Biomass Burning (as light slowly improves on the descending passes, conditions are ideal to document the more subtle smoke loadings in the atmosphere over this most flammable part of the globe.  Looking left for source fires in the dense Angolan savanna; looking right for the views of this mass moving out into the Atlantic Ocean.  The latter views–of the haze “exhaust” from the continent–are being specifically requested by scientists), and Marietta, Ohio (LEWIS & CLARK SITE(S):  Marietta is noted as the oldest settlement on the Ohio River.  It is located about ten miles upriver from Parkersburg, West Virginia, where the Muskingum River enters from the north and the Ohio makes a large meander north [looking just right of track].  In the summer of 1803, on their return to St. Louis, Lewis and Clark stopped here for a visit.  It was one of the few settlements they encountered on their way west).
CEO images can be viewed at the websites

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 12:40pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On (18 amp mode).  Vozdukh CO2 scrubber is On (in Manual Mode 3/5).  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  MCA (major constituents analyzer) is operating.  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On; SKV-2 is Off.
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 28.3; ppO2 (mmHg) — data invalid; ppCO2 (mmHg) — data invalid.
  • SM Transfer Compartment:  Pressure (mmHg) — 751; temperature (deg C) — 20.9.
  • FGB Cabin:  Pressure (mmHg) — 752; temperature (deg C) — 22.0.
  • Node:  Pressure (mmHg) — 745.80; temperature (deg C) — 24.5 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 747.24; temperature (deg C) — 23.0; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):  Pressure (mmHg) — 747.34; temperature (deg C) — 30.6; shell heater temp (deg C) — 25.4, ppO2 (mmHg) — 172.1; ppCO2 (mmHg) — 4.1.
  • PMA-1:  Shell heater temp (deg C) — 25.2.
  • PMA-2:  Shell heater temp (deg C) — 14.0.

(n/a = data not available)

Propulsion System (PS):

  • Total propellant load available: 3803 kg (8384 lb) as of 6/19  [SM(774) + FGB(2447) + Progress M(182) +Progress M-1(400)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in AutoTrack.
  • SM batteries:  Battery #4 is disconnected (failed 6/16); all other batteries (7) are in “Partial Charge” mode.
  • FGB batteries:  Battery #4 is in “Cycle” mode; all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-1 is operating; INT-2 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-2 MDM is Off; PL-1 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • XPOP (x-axis perpendicular to orbit plane = “sun-fixed” [yaw: 0.5 deg, pitch: -9.0 deg., roll: 0 deg]), with CMG TA (thruster assist).Management.
  • Solar Beta angle:  -25.6 deg (magnitude decreasing).

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally [IAC-1 (internal audio controller #1) being analyzed after self-test error.  IAC-2 is prime.]
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at MBS PDGF #1 with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 7:29am EDT [= epoch]):

  • Mean altitude — 388.1 km
  • Apogee — 392.7 km
  • Perigee — 383.4km
  • Period — 92.32 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.0006839
  • Orbits per 24-hr. day — 15.60
  • Solar Beta Angle — -25.6 deg (magnitude decreasing)
  • Mean altitude loss in last 24 hours — 70 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 26214
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.